Flexible Budget for Assembly Department Cabinaire Inc. is one of the largest manufacturers of office furniture...

  1. Flexible Budget for Assembly Department

    Cabinaire Inc. is one of the largest manufacturers of office furniture in the United States. In Grand Rapids, Michigan, it assembles filing cabinets in an Assembly Department. Assume the following information for the Assembly Department:

    Direct labor per filing cabinet 20 minutes
    Supervisor salaries $147,000 per month
    Depreciation $21,000 per month
    Direct labor rate $15 per hour

    Prepare a flexible budget for 11,000, 14,000, and 17,000 filing cabinets for the month of March in the Assembly Department similar to Exhibit 5. Enter all amounts as positive numbers.

    Cabinaire Inc.
    Assembly Department Budget
    Month Ending March 31 (assumed data)
    Units of production 11,000 14,000 17,000
    Variable cost:
    • Depreciation
    • Direct labor
    • Supervisor salaries
    • Rent
    • Utilities
    $ $ $
    Total variable cost $ $ $
    Fixed cost:
    • Direct labor
    • Electric power
    • Rent
    • Supervisor salaries
    • Wages
    $ $ $
    • Depreciation
    • Direct labor
    • Insurance
    • Utilities
    • Wages
    Total fixed cost $ $ $
    Total department costs $ $ $

    Feedback

Loading item

There was an error loading this item. If this continues to occur, please contact Technical Support.

Check My Work

  • Previous
  • Next
  • 100% Correct
  • Partially Correct
  • Incorrect
  • Needs Instructor Grading

Basic Calculatorclose

0

UseEntBSBSpCEHomCEnd

789+

456-

123*

0.=/
Answer #1
Cabinaire Inc.
Assembly Department Budget
Month Ending March 31 (assumed data)
Units of production 11000 14000 17000
Variable cost:
Direct labor 55000 70000 85000
Total variable cost 55000 70000 85000
Fixed cost:
Supervisor salaries 147000 147000 147000
Depreciation 21000 21000 21000
Total fixed cost 168000 168000 168000
Total department costs 223000 238000 253000
Workings:
Units of production Direct labor
11000 55000 =11000*15*(20/60)
14000 70000 =14000*15*(20/60)
17000 85000 =17000*15*(20/60)